OpenStudy (anonymous):

Find three consecutive odd integers such that the product of the first and the second exceeds the third by 8

OpenStudy (anonymous):

The first two integers can be called x and x + 2 What would we call the third one?

OpenStudy (anonymous):

{n - 2, n, n + 2} ((n - 2) n) = (n + 2) + 8 n=5 {3, 5, 7}

OpenStudy (anonymous):

i would guess and check for this one

OpenStudy (anonymous):

Your teacher gives credit for guess and check?